### v1

# Spatial Context-Driven Positive Pair Sampling for Enhanced Histopathology Image Classification

Abstract
Deep learning has demonstrated great promise in cancer classification from whole-slide images (WSIs) but remains constrained by the need for extensive annotations. Annotation-free methods, such as multiple instance learning (MIL) and self-supervised learning (SSL), have emerged to address this challenge; however, current SSL techniques often depend on synthetic augmentations or temporal context, which may not adequately capture the intricate spatial relationships inherent to histopathology. In this work, we introduce a novel spatial context-driven positive pair sampling strategy for SSL that leverages the natural coherence of adjacent patches in WSIs. By constructing biologically relevant positive pairs from spatially proximate patches, our approach harnesses inherent spatial coherence to enhance patch-level representations, ultimately boosting slide-level classification performance. Experiments on multiple datasets reveal that our strategy improves classification accuracy by 5% to 10% over the standard method, paving the way for more clinically relevant AI models in cancer diagnosis. The code is available at https://anonymous.4open.science/r/contextual-pairs-E72F/.
Keywords:
Cancer Diagnosis Digital Pathology Whole-Slide Images Self-Supervised Learning Positive Pair Sampling.
1 Introduction
Deep learning (DL) has transformed cancer classification with whole-slide images (WSIs), achieving pathologist-level performance [15, 19, 20]. However, training these models requires large labeled datasets, which are scarce and labor-intensive to annotate, as they depend on expert pathologists [17, 25].
To address the scarcity of labeled data, annotation-free approaches have been proposed [14, 16, 23, 26]. One prominent method is Multiple Instance Learning (MIL), which treats each WSI as a "bag" of image patches with a single slide-level label [7]. Another promising strategy is self-supervised learning (SSL), which enables models to learn from vast amounts of unlabeled data. SSL has proven particularly valuable in histopathology, where obtaining expert-annotated datasets is costly and time-consuming. Notably, SSL pre-training on pathology data consistently outperforms ImageNet pre-training, leading to better representations for downstream pathology tasks [15, 19, 20].
While SSL enables feature learning without labeled data [15, 18], adapting it to pathology images presents unique challenges. Unlike natural images, pathology images lack a canonical orientation and exhibit low color variation [15]. Standard SSL techniques, such as random flips, stain augmentations, and multi-view training, attempt to address these issues [8, 24, 6, 15]. However, these synthetic transformations may not fully capture the natural variations found in real-world tissue samples.
Recent work in natural image domains suggests that leveraging spatial or temporal relationships within unlabeled data can generate more meaningful positive pairs for SSL [21, 13]. For example, Pantazis et al. [21] utilized spatial and temporal proximity to create informative training pairs in wildlife camera images, while Tile2Vec [13] employed geographic proximity with a triplet loss to differentiate between similar and dissimilar regions in satellite images. Inspired by these findings, we propose a novel SSL approach that incorporates spatial proximity for more biologically relevant positive pair sampling in WSI analysis.
Figure 1: In the standard approach (blue box), a patch from the WSI is transformed to create a standard positive pair for self-supervised learning. Our method (green box) additionally leverages spatial proximity to sample a near-neighbor patch, forming a contextual positive pair. By combining both pairs, we exploit transformation invariance and contextual similarity, resulting in more robust representation learning.
In this paper, we introduce and evaluate our contextual positive pair sampling strategy within the SSL framework (illustrated in Figure 1). Rather than aiming to surpass state-of-the-art (SOTA) performance, our goal is to present a biologically grounded method that enhances representation learning for histopathology, aligning with the growing interest in foundational models for medical imaging [15, 20, 19]. By leveraging spatial coherence between adjacent image patches, our approach constructs more meaningful positive pairs, leading to improved WSI classification performance over standard SSL sampling strategies.
2 Methodology
2.1 Self-supervised Learning
Self-supervised Learning (SSL) methods aim to train an embedding function that maps an input image $x$ to a meaningful representation, enabling downstream tasks without requiring extensive labeled data. Recent SSL methods adopt a joint embedding architecture, where two networks are optimized to produce similar embeddings for different augmented views of the same image [3, 10, 28, 1, 5].
In the context of WSI analysis, an input image $x$ is a small patch extracted from a WSI, denoted as $P=\{p_{1},...,p_{m}\}$. Given a set of image transformation $T$, transformations $t_{1},t_{2}\sim T$ are sampled to two views: $v_{1,i}=t_{1}(p_{i})$ and $v_{2,i}=t_{2}(p_{i})$. An encoder $f$ is used to produce representations $z_{1}=f(v_{1,i})$ and $z_{2}=f(v_{2,i})$. A projector $h$ is commonly used to produce projections $y_{1}=h(z_{1})$ and $y_{2}=h(z_{2})$. A loss function $L(y_{1},y_{2})$ is then minimized to enforce similarity in the embedding space. Subsequently, the learned patch-level embeddings are aggregated into a slide-level representation through a features aggregator. Finally, this slide-level representation is utilized by a dedicated slide-level model for downstream classification tasks, such as cancer diagnosis.
Although earlier SSL methods like SimCLR [6] rely on negative pairs, recent approaches opt to not use negative pairs primarily for computational efficiency and improved scalability [10, 28, 1, 5]. Moreover, in benign WSIs, all extracted patches originate from benign tissue, making it infeasible to construct true negative pairs. Consequently, our method focuses exclusively on positive pairs.
2.2 Contextual Positive Pairs Sampling
The overview of our method is illustrated in Fig 1. Our framework leverages spatial proximity to construct positive pairs that capture shared underlying characteristics. Before training the SSL model, we sample $M$ patches, termed pivot patches, from all WSIs to form the training set $U^{P}$. For each pivot patch, we identify $K$ spatially proximate patches within the same WSI, referred to as neighbor patches ($U^{N}$), based on the Chebyshev distance [4] (given that WSIs are fundamentally grid-based structures). During training, a neighbor patch is randomly selected for each pivot patch, provided its distance falls within predefined thresholds, forming a contextual pair $\{p^{i}_{j}\in U^{P},p^{i}_{k}\in U^{N}\}$. In addition, we construct standard positive pairs for each pivot patch by applying two independent transformations, $t_{1}$ and $t_{2}$, yielding $\{t_{1}(p^{i}_{j}),t_{2}(p^{i}_{j})\}$. The SSL loss is computed using two components: one based on the contextual pair $(v_{1,j}=t_{1}(p^{i}_{j}),v_{1,k}=t_{1}(p^{i}_{k}))$ and another based on the standard augmentation-based pair $(v_{1,j}=t_{1}(p^{i}_{j}),v_{2,j}=t_{2}(p^{i}_{j}))$. These loss terms are weighted by a hyperparameter $\alpha$, as formulated in Equation 1:
$L^{\prime}=\alpha\cdot L(v_{1,j},v_{1,k})+(1-\alpha)\cdot L(v_{1,j},v_{2,j})$ (1)
This formulation balances contextual similarity and transformation invariance, guiding the model to learn robust representations. Moreover, it is agnostic to the specific SSL method employed, making it adaptable to various frameworks. Moreover, while various distance metrics exist, we selected Chebyshev distance because WSIs are fundamentally grid-based structures, with patches arranged in a matrix-like format.
3 Experiments and Results
3.1 Datasets
3.1.1 Public Datasets
We evaluated our method on two publicly available datasets: Camelyon16 [2] and TCGA-NSCLC. The Camelyon16 (Breast) dataset consists of 399 WSIs of lymph node sections, including 257 normal slides and 142 slides with metastatic regions. It is divided into a training set of 270 slides and a test set of 129 slides. We further partitioned the training set into 80 normal and 80 malignant slides for training, with an additional 80 normal and 30 malignant slides reserved for validation. TCGA-NSCLC (Lung) dataset comprises 1,046 WSIs, with 836 slides used for training and 210 for testing, spanning two lung cancer subtypes: 534 Lung Adenocarcinoma (LUAD) and 512 Lung Squamous Cell Carcinoma (LUSC). On both datasets, we extracted non-overlapping 256x256 pixel tiles from the tissue regions at 20x magnification.
3.1.2 Private Datasets
We evaluated our method in two private datasets acquired from a medical foundation: (1) 600 stomach WSIs and (2) 600 colorectal WSIs. Each WSI contains 2 to 4 slices from the same tissue sample. Among the slides, and for each dataset, 200 were classified as normal/benign slides (N), 200 were classified as tissues in dysplasia (D), and 200 were classified as cancerous/malignant (M). The Institutional Review Board (Approval #SMF-IRB-2020–007) of the organization and the Institutional Review Board (Approval #KAIST-IRB-22-334, #KH2020-116) of Korea Advanced Institute of Science and Technology (KAIST) have both granted approval for research use. The WSIs were stained with hematoxylin and eosin (H&E), digitized as MIRAX files with a sensor resolution of 200x, using a Panoramic Flash 250 III scanner, and were examined and graded by two pathologists from the same organization. All patient records were anonymized. Moreover, for each class, we randomly split the WSIs into 140 slides for training, 30 slides for validation, and 30 slides for testing. Using a sliding window approach, we extracted 256x256 pixel patches at 20x magnification with the OpenSlide Python library [9].
3.2 Implementation Details
We conducted our experiments using PyTorch [22] on a Linux workstation with a single 32 GB NVIDIA GeForce RTX 3080 Ti GPU. We employed three SSL methods—Barlow Twins [28], BYOL [10], and VICReg [1]—to train the patch-level representation encoders. ResNet18 [11] was used as the backbone architecture for the SSL methods. The optimization was performed using the LARS optimizer [27], configured with a weight decay of $1\times 10^{-6}$. Separate learning rates were set for biases and weights at 0.0048 and 0.2, respectively. A batch size of 1024 was employed, and all models were trained for 100 epochs.
To generate the augmented views, we used a diverse set of transformations inspired by [6] and [15]. These augmentations included vertical and horizontal flipping, random resized cropping, color jittering, and Gaussian blur, which helped the model become invariant to spatial and visual variations.
We aggregated the patch features using Multiple Instance Learning (MIL) to make a slide-level prediction. Specifically, we used attention-based MIL (ABMIL) [12]. During MIL training, we employed the Adam optimizer with a weight decay of $1\times 10^{-2}$, a learning rate of 0.001, and a total of 100 epochs. The performance of the models was evaluated using accuracy and the Area Under the Receiver Operating Characteristic Curve (AUROC).
3.3 Contextual Positive Pair Sampling
Mining neighboring patches based on distance in WSIs carries the risk of sampling false positive pairs, referred to as mismatches. To quantify this risk, we calculate the percentage of mismatches as the sampling distance increases, as illustrated in Figure 2. Note that we calculated mismatching percentages only for WSIs annotated as non-benign (malignant for Camelyon16; Dyplasia and malignant for the private datasets), as benign WSIs contain only benign patches. Additionally, we excluded the TCGA dataset due to the absence of patch-level labels.
Figure 2: Percentage of mismatches as a function of Chebyshev distance for three histopathology datasets (Stomach, Colon, and Breast).
As the distance grows, the likelihood of mismatches rises, reflecting that patches drawn from more distant regions are more likely to belong to different classes or tissue types. Based on this finding, we sampled patches adjacent to the pivot patches as positive pairs, enforcing a maximum Chebyshev distance of 1. During training, neighbor patches are selected randomly. Moreover, we set $\alpha=0.5$ to maintain an equal balance between contextual similarity and transformation invariance. The impact of distance constraints and $\alpha$ was further examined in our ablation studies. Additionally, we explored a scenario where neighbor patches were sampled randomly without any distance constraints.
3.4 Slide-level Classification
Table 1 compares standard positive pair sampling with two contextual variants: Context($\infty$), which samples neighbor patches without distance constraints, and Context($1$), which restricts sampling to patches within a Chebyshev distance of 1. Across all datasets and SSL methods, Context($1$) consistently outperforms the standard approach in both accuracy and AUROC, highlighting the effectiveness of leveraging immediate spatial context.
Table 1: Slide-level performance across datasets, reporting accuracy and AUROC for Positive Pair Sampling strategies. Gray cells indicate contextual approaches: ($\infty$) for unrestricted distance sampling and ($1$) for Chebyshev distance of $1$. BT denotes Barlow Twins. Bold indicates highest value. Results are averaged over five runs.
Dataset SSL Accuracy AUROC
Standard Context($\infty$) Context($1$) Standard Context($\infty$) Context($1$)
Private (Stomach) BT 0.827 0.836 0.893 0.941 0.947 0.966
BYOL 0.813 0.833 0.902 0.954 0.964 0.972
VICReg 0.831 0.816 0.893 0.947 0.945 0.960
Private (Colon) BT 0.820 0.844 0.891 0.926 0.945 0.971
BYOL 0.802 0.756 0.916 0.922 0.930 0.987
VICReg 0.807 0.809 0.916 0.926 0.931 0.981
Camelyon16 (Breast) BT 0.710 0.710 0.763 0.656 0.658 0.722
BYOL 0.757 0.767 0.798 0.751 0.738 0.753
VICReg 0.757 0.716 0.798 0.726 0.705 0.777
TCGA (Lung) BT 0.773 0.771 0.825 0.839 0.835 0.909
BYOL 0.763 0.760 0.849 0.847 0.844 0.920
VICReg 0.770 0.769 0.857 0.846 0.846 0.914
The performance of Context($\infty$) is more variable—it sometimes matches or exceeds the standard approach but occasionally underperforms, suggesting that unconstrained sampling may introduce noise or less relevant pairs. On the other hand, the magnitude of improvement with Context($1$) varies across datasets. For instance, the Colon dataset sees accuracy gains exceeding 10% (from 0.807 to 0.916 with VICReg), while the Breast dataset improves by about 4% (from 0.757 to 0.798 with VICReg). This variation likely reflects differences in dataset characteristics, such as the presence of multiple tissue samples per WSI or inherent structural complexity. Despite these variations, the consistent advantage of Context($1$) across all datasets underscores the importance of spatially coherent sampling for learning more robust histopathology representations. These findings suggest that leveraging spatial context in SSL not only improves performance in slide-level classification but also provides a biologically intuitive framework for annotation-free learning in digital pathology.
3.5 Ablation Studies
3.5.1 Proportion of Contextual Similarity
Our loss formulation balances contextual similarity and transformation invariance, weighted by $\alpha$. Figure 3 shows the effect of such proportion on the classification accuracy for all the datasets. The data indicates that incorporating the contextual information ($\alpha>0$) yields accuracy gains relative to a baseline ($\alpha=0$). Moreover, while $\alpha=1$ (only neighbor-based pairs) outperforms $\alpha=0$, it is rarely the absolute best. In most datasets and methods, a mixture of contextual and standard augmentation-based pairs yields the highest accuracy. However, the magnitude and behavior of these gains vary by both method and dataset. For instance, BYOL shows more variable performance, with sensitivity to the $\alpha$ value, while VICReg generally provides consistent improvements. Overall, these results validate the key idea that balancing “transformation invariance” and “contextual similarity” is beneficial. The exact sweet spot for $\alpha$ varies, but typically falls in the middle range, underscoring the importance of simultaneously leveraging both types of positive pairs in self-supervised histopathology image analysis.
Figure 3: Accuracy gain relative to the baseline (without contextual information; i.e., $\alpha=0$) for each SSL method across four datasets. Results are averaged over five runs.
3.5.2 Effect of Neighboring Distance
Figure 4: Effect of neighboring distance on slide-level classification accuracy measured via accuracy gain (%). Error regions represent the standard deviation.
We investigated how distance thresholds for selecting contextual positive pairs affect SSL performance in slide-level classification. We sampled 100,000 pivot patches and identified neighbor patches within specific distance thresholds, covering both short- and long-range contexts.
Figure 4 illustrates how neighboring distance affects classification accuracy across all datasets. Overall, shorter distances yield higher accuracy gains, while performance typically declines when the distance exceeds 5—likely due to diminished contextual relevance and increased noise or mismatches. Although all three SSL methods benefit from closer patch sampling, the magnitude of improvement varies with dataset-specific factors such as tissue heterogeneity and the SSL method employed. Interestingly, accuracy in the Stomach and Colon datasets rebounds at greater distances, a trend we attribute to these WSIs containing multiple tissue samples. Based on these observations and our false positive pair analysis, we recommend sampling neighbor patches immediately adjacent to the pivot (i.e., $distance=1$).
4 Discussion
We introduced a spatial context-driven positive pair sampling strategy that enhances self-supervised learning (SSL) in whole-slide image (WSI) analysis by leveraging the inherent spatial coherence of tissue morphology. Our experimental results show that integrating contextual pairs improves slide-level classification under Multiple Instance Learning, with ablation studies indicating that smaller neighboring distances generally yield greater gains—although the optimal distance varies by dataset. Excessive reliance on spatial proximity alone can reduce representation quality, highlighting the need to balance contextual and standard augmentations. While our study did not aim to surpass state-of-the-art performance benchmarks, we focused on establishing a biologically meaningful SSL framework that improves representation learning in computational histopathology. Our findings underscore the importance of spatial context in guiding SSL models toward more diagnostically relevant feature extraction, which could be particularly beneficial for training Pathology Foundation Models on large-scale, annotation-limited datasets. Additionally, optimizing our approach for computational efficiency and exploring its applicability across diverse tissue types and staining variations remain key areas for future research. In conclusion, by aligning positive pair formation with the inherent structure of WSIs, our method provides a promising direction for self-supervised histopathology analysis, paving the way for more scalable and generalizable AI-driven pathology models.