### v2

# Leveraging Spatial Context for Positive Pair Sampling in Histopathology Image Representation Learning

Abstract
Deep learning has shown strong potential in cancer classification from whole-slide images (WSIs), but the need for extensive expert annotations often limits its success. Annotation-free approaches, such as multiple instance learning (MIL) and self-supervised learning (SSL), have emerged as promising alternatives to traditional annotation-based methods. However, conventional SSL methods typically rely on synthetic data augmentations, which may fail to capture the spatial structure critical to histopathology. In this work, we propose a spatial context-driven positive pair sampling strategy that enhances SSL by leveraging the morphological coherence of spatially adjacent patches within WSIs. Our method is modular and compatible with established joint embedding SSL frameworks, including Barlow Twins, BYOL, VICReg, and DINOv2. We evaluate its effectiveness on both slide-level classification using MIL and patch-level linear probing. Experiments across four datasets demonstrate consistent performance improvements, with accuracy gains of 5% to 10% compared to standard augmentation-based sampling. These findings highlight the value of spatial context in improving representation learning for computational pathology and provide a biologically meaningful enhancement for pretraining models in annotation-limited settings. The code is available at https://anonymous.4open.science/r/contextual-pairs-E72F/.
1 Introduction
Deep learning (DL) has revolutionized cancer diagnosis using whole-slide images (WSIs), achieving pathologist-level performance in several tasks [21, 25, 27]. However, these advances rely heavily on large, expert-annotated datasets, which is an expensive and time-consuming requirement that limits scalability across institutions and cancer types [23, 41].
To mitigate the annotation bottleneck, label-efficient approaches such as Multiple Instance Learning (MIL) [9] and self-supervised learning (SSL) [19, 22, 33, 43] have gained traction. In particular, SSL has emerged as a powerful tool for learning visual representations directly from unlabeled WSI data. SSL pretraining on histopathology data has consistently outperformed ImageNet pretraining, yielding improved features for downstream classification and segmentation tasks [21, 25, 27].
Despite its promise, applying SSL to pathology introduces unique challenges. WSIs differ from natural images in that they lack consistent orientation, exhibit low color variability, and contain complex spatial tissue structures [21]. Common SSL strategies (e.g., synthetic transformations like flips or stain augmentations [10, 35, 8, 21]) may not fully capture the biological variability inherent in tissue morphology.
Recent advances in natural image domains highlight the benefit of using contextual signals, spatial or temporal, to define semantically meaningful positive pairs [30, 18]. For instance, Tile2Vec [18] uses geographic proximity in satellite imagery with a triplet loss, while Pantazis et al. leverage spatiotemporal co-occurrence in wildlife monitoring. However, these methods are not directly applicable to histopathology, where (i) extremely high-resolution WSIs form grid-like spatial layouts, (ii) semantic class transitions can be subtle, and (iii) true negatives are difficult to define due to localized disease features.
Figure 1: In standard SSL (blue), a patch is transformed to generate a positive pair. Our method (green) augments this by incorporating a spatially adjacent neighbor as a contextual pair. This leverages biological structure to enforce both transformation invariance and spatial coherence in representation learning.
In this paper, we propose a spatially coherent positive pair sampling strategy tailored for histopathology SSL. Rather than relying solely on augmentations, we incorporate the natural proximity of WSI patches to generate biologically meaningful positive pairs (Figure 1). Our approach is method-agnostic and can be integrated into contrastive or redundancy-reduction SSL frameworks.
We evaluate our method across multiple datasets and architectures. Rather than aiming to surpass state-of-the-art performance, our goal is to present a biologically grounded method that enhances representation learning for histopathology, aligning with the growing interest in foundational models for medical imaging [21, 27, 25]. By leveraging spatial coherence between adjacent image patches, our approach constructs more meaningful positive pairs, leading to improved WSI classification performance over standard SSL sampling strategies.
2 Related Work
2.1 Deep Learning for Digital Histopathology
Deep learning (DL) approaches for whole-slide image (WSI) classification typically follow a multi-stage pipeline [36, 31, 3, 34]. A WSI is first segmented into smaller patches, often with non-tissue regions removed. These patches are then processed (either labeled at the patch level or treated as unlabeled instances) and their features aggregated into slide-level representations using attention-based pooling [9, 40, 44], feature maps [36, 31], or graphs [47, 3]. Finally, a downstream model performs slide-level classification.
However, despite the abundance of WSIs, two challenges limit supervised DL in pathology. First, generating high-quality annotations is time-intensive and requires expert pathologists, with risks of fatigue and inter-observer variability [6, 2]. Second, the gigapixel scale of WSIs makes end-to-end training computationally prohibitive. Furthermore, class imbalance and rare diseases exacerbate data scarcity for specific tasks [17, 12].
To overcome these barriers, several annotation-efficient approaches have been explored: transfer learning [42, 1, 37], synthetic data augmentation [20, 11, 39, 38], and annotation-free learning [19, 22, 33]. The latter, particularly self-supervised learning (SSL), has shown promise in extracting meaningful features without pixel-level labels.
2.2 Self-supervised Learning in Histopathology
Self-supervised learning (SSL) aims to learn visual representations from unlabeled data by solving proxy tasks. In histopathology, patch-level embeddings learned through SSL can be aggregated via multiple instance learning (MIL) to perform downstream WSI classification [44]. This paradigm reduces reliance on curated patch labels while achieving competitive performance.
Contrastive learning is a widely used SSL approach where a model learns to align representations of transformed views of the same input (positive pairs) while pushing apart other inputs (negatives) [8]. However, in WSIs, defining negatives is nontrivial: benign tissue dominates most slides, and spatially distant patches may still share morphology. Recent SSL frameworks such as BYOL, Barlow Twins, VICReg, and DINO avoid the need for negative pairs altogether [14, 46, 4, 7], focusing on redundancy reduction or view alignment instead.
Still, most SSL approaches rely heavily on data augmentations (e.g., cropping, color jitter, flipping) designed for natural images. These transformations may not adequately capture biological variations in tissue, especially structural continuity or local morphological patterns.
Foundation models for histopathology have recently emerged as a promising direction to scale representation learning across diverse WSI datasets and clinical tasks [27, 25, 21]. These models typically use large transformer-based architectures and are trained with extensive unlabeled data using objectives such as masked image modeling or contrastive learning. While this study is not designed to compete with or replace these large-scale models, it introduces a biologically motivated sampling strategy that can be integrated into existing SSL frameworks with minimal overhead. The incorporation of spatial context offers a way to improve patch-level representation quality, which could serve as a complementary enhancement to pretraining routines, especially in scenarios where data efficiency or morphological relevance is critical.
2.3 Positive Pair Sampling and Spatial Context
Recent work has shown that the choice of positive pairs can significantly affect SSL performance; sometimes more than the backbone architecture or loss function [30, 18, 24]. For example, Pantazis et al. leveraged spatial and temporal cues in wildlife camera imagery to form semantically meaningful pairs. Tile2Vec [18] introduced spatial sampling in satellite imagery, using geographic proximity and triplet loss to learn local structure-aware embeddings.
These works suggest that natural context (rather than synthetic augmentation) can provide rich supervisory signals. However, they have not been adapted to the unique structure of WSIs, which are high-resolution, grid-structured, and label-scarce.
In this work, we explore positive pair sampling based on spatial proximity within WSIs. Adjacent patches in a WSI typically exhibit strong morphological coherence, offering a biologically grounded signal for SSL. Unlike Tile2Vec, we do not assume distant patches are dissimilar: many WSIs, especially normal ones, contain spatially repetitive benign tissue. Also, we do not use a triplet loss or require negatives, aligning our approach with modern SSL practices.
Similar to our work is SimTriplet: both approaches exploit the spatial continuity of tissue morphology in WSIs to define biologically meaningful supervisory signals for self-supervised learning. Specifically, both methods construct positive pairs from spatially adjacent patches under the assumption that proximity implies semantic similarity in histopathology. However, our approach differs from SimTriplet in both architecture design and training objective. SimTriplet employs a triplet-based network that explicitly optimizes intra-sample and inter-sample similarities using a custom triplet structure and loss. In contrast, our method introduces a flexible contextual positive pair sampling strategy that does not require architectural changes or triplet loss formulations. Instead, we augment standard SSL pipelines by blending augmentation-based and context-based positive pairs through a weighted combination in the loss function (Equation 1). This modular design allows seamless integration into popular joint embedding SSL frameworks (e.g., BYOL, Barlow Twins, VICReg, DINOv2) without relying on handcrafted triplet structures or predictors. Additionally, we explicitly investigate the impact of spatial distance in contextual sampling, providing deeper insight into how spatial granularity influences representation learning in WSI data.
3 Methodology
3.1 Overview
We present a self-supervised learning (SSL) framework for whole-slide image (WSI) analysis that incorporates spatial context into positive pair selection. The pipeline consists of three main components: (1) patch extraction from WSIs, (2) SSL pretraining using contextual positive pair sampling, and (3) downstream evaluation via multiple instance learning (MIL) and linear probing. Our method is modular and compatible with a variety of joint embedding SSL approaches.
3.2 Contextual Positive Pair Sampling
We introduce a spatially-aware positive pair sampling strategy that leverages the local structural coherence of tissue morphology. The key idea is to incorporate spatially adjacent patches as contextual signals, augmenting standard augmentation-based sampling. An overview of this strategy is shown in Figure 1.
Let $U^{P}$ denote a set of $M$ pivot patches sampled from the WSIs in the training set. For each pivot patch $p^{i}_{j}\in U^{P}$, we identify a set of $K$ neighbor patches $U^{N}=\{p^{i}_{k}\}$ within the same WSI, constrained by a maximum Chebyshev distance. This metric is chosen due to its alignment with the grid structure of WSIs, where patches are spatially indexed.
During training, we construct two types of positive pairs:
  1. 1.
    Standard augmentation-based pair: Two augmented views of the same patch, $v_{1,j}=t_{1}(p^{i}_{j})$ and $v_{2,j}=t_{2}(p^{i}_{j})$, where $t_{1},t_{2}\sim T$ are sampled transformations.
  2. 2.
    Contextual pair: A neighbor patch $p^{i}_{k}\in U^{N}$ is randomly selected, and both $p^{i}_{j}$ and $p^{i}_{k}$ are augmented using the same transformation $t_{1}$, producing $v_{1,j}=t_{1}(p^{i}_{j})$ and $v_{1,k}=t_{1}(p^{i}_{k})$.
The combined loss function is defined as:
$L^{\prime}=\alpha\cdot L(v_{1,j},v_{1,k})+(1-\alpha)\cdot L(v_{1,j},v_{2,j})$ (1)
where $L(\cdot,\cdot)$ denotes the loss function from the SSL framework in use (e.g., redundancy reduction for Barlow Twins or BYOL). The weight $\alpha\in[0,1]$ balances contextual similarity and transformation invariance.
This formulation avoids the need for negative pairs and is particularly well-suited for WSI data, where distant patches may still belong to the same class. The method is biologically motivated, computationally efficient, and compatible with both contrastive and non-contrastive SSL paradigms.
3.3 Integration with SSL Frameworks
Our sampling strategy is architecture-agnostic and integrates seamlessly with joint embedding SSL frameworks such as Barlow Twins, BYOL, VICReg, and DINOv2. During training, contextual and standard positive pairs are combined as defined in Equation 1, guiding the encoder to learn representations that are robust to both augmentations and local spatial variations. This design allows the approach to serve as a plug-and-play module within existing WSI pipelines, improving patch-level feature quality with minimal overhead.
4 Experiments and Results
4.1 Datasets
Public Datasets
We evaluated our method on two publicly available datasets: Camelyon16 [5] and TCGA-NSCLC. The Camelyon16 (Breast) dataset consists of 399 WSIs of lymph node sections, including 257 normal slides and 142 slides with metastatic regions. It is divided into a training set of 270 slides and a test set of 129 slides. We further partitioned the training set into 80 normal and 80 malignant slides for training, with an additional 80 normal and 30 malignant slides reserved for validation. TCGA-NSCLC (Lung) dataset comprises 1,046 WSIs, with 836 slides used for training and 210 for testing, spanning two lung cancer subtypes: 534 Lung Adenocarcinoma (LUAD) and 512 Lung Squamous Cell Carcinoma (LUSC). On both datasets, we extracted non-overlapping 256x256 pixel tiles from the tissue regions at 20x magnification.
Private Datasets
We evaluated our method on two private datasets obtained from a collaborating medical institution: (1) 600 whole slide images (WSIs) of stomach tissue and (2) 600 WSIs of colorectal tissue. Each WSI comprises 2 to 4 sections from the same specimen. For each dataset, the slides were evenly categorized into normal/benign (N), dysplastic (D), and malignant (M) groups, with 200 slides per class. For each class, we randomly split the slides into 140 for training, 30 for validation, and 30 for testing. Patches of size 256$\times$256 pixels were extracted at 20$\times$ magnification using a sliding window approach implemented with the OpenSlide Python library [13].
Research use of these datasets was approved by the Institutional Review Boards of Institution A (Approval #InstA-IRB-2020–007) and Institution B (Approval #InstB-IRB-22-334, #InstB-2020-116). All WSIs were stained with hematoxylin and eosin (H&E), digitized in MIRAX format at a resolution of 200x using a Panoramic Flash 250 III scanner, and independently reviewed by two board-certified pathologists affiliated with Institution A. Patient data were fully anonymized prior to analysis.
4.2 Implementation Details
All experiments were conducted using PyTorch [32] on a Linux workstation equipped with a single NVIDIA GeForce RTX 3080 Ti GPU (32 GB). We evaluated four self-supervised learning (SSL) methods—Barlow Twins [46], BYOL [14], VICReg [4], and DINOv2 [29]—for training patch-level representation encoders. All methods used ResNet-18 [15] as the backbone, except DINOv2, which employed a Vision Transformer (ViT-Tiny) architecture.
For Barlow Twins, BYOL, and VICReg, we used the LARS optimizer [45] with a weight decay of $1\times 10^{-6}$. The learning rates were set to 0.2 for weights and 0.0048 for biases. These models were trained for 100 epochs with a batch size of 1024. For DINOv2, we adopted the AdamW optimizer with a cosine learning rate scheduler. The initial learning rate was set to 0.0005 with a weight decay of 0.04. The teacher’s temperature was linearly warmed up from 0.02 to 0.04 over 15 epochs. Due to computational constraints, DINOv2 was trained for 50 epochs using a batch size of 128.
Image Augmentation
To generate augmented views for self-supervised learning, we employed a diverse set of image transformations inspired by SimCLR [8], Kang et al. [21], and DINOv2 [29]. The augmentations included random horizontal and vertical flipping, random resized cropping, color jittering, and Gaussian blur. Additionally, for DINOv2, we adopted a multi-crop strategy consisting of two global crops and eight local crops. The global crops were generated using random resized cropping (scale range of 0.5–1.0) with bicubic interpolation, while the local crops were produced with smaller resized crop regions (scale range of 0.05–0.4). All crops were normalized using standard ImageNet mean and standard deviation statistics.
Slide-level Classification
To perform slide-level classification, we aggregated patch-level features using a Multiple Instance Learning (MIL) framework. In this setting, each WSI was treated as a bag of instances (i.e., extracted patches), and a single label was assigned to the entire bag rather than to individual patches. Specifically, we adopted the attention-based MIL (ABMIL) approach proposed in [16], which learns to assign attention weights to each patch feature. For each WSI, we first extracted patch embeddings using the pretrained encoders without any fine-tuning. These patch embeddings were then input to the ABMIL model to predict the slide-level label. The model was trained using the Adam optimizer with a learning rate of 0.001 and weight decay of $1\times 10^{-2}$. Training was conducted for 100 epochs.
Linear Probing
To evaluate the quality of the learned representations, we performed linear probing by training a shallow classifier on top of the frozen encoders obtained from the SSL methods. We sampled 5,000 patches per class for training and 2,500 patches per class for both validation and testing. The encoder weights were loaded from the corresponding SSL-pretrained checkpoints, and all encoder parameters were frozen during training. We evaluated both ResNet-18 and ViT-Tiny backbones. For ResNet-18, we used an SGD optimizer with an initial learning rate of 0.1 and applied cosine annealing as the learning rate scheduler. The linear classifier was trained for 50 epochs. For ViT-Tiny, we used the AdamW optimizer with a learning rate of 3e-4 and weight decay of 0.01, combined with a cosine scheduler and 5 warm-up steps over 50 training epochs. In both setups, the frozen encoder outputs were fed into a linear layer to perform tissue classification.
As a supervised baseline, we trained the same ResNet-18 and ViT-Tiny backbones from scratch using labeled patch data. The training configuration, including optimizers, learning rates, and schedulers, was identical to that used in the linear probing setup. However, unlike linear probing, the entire model (encoder and classifier) was trained end-to-end for 50 epochs.
We excluded the TCGA dataset from the linear probing and supervised training experiments, as it does not contain patch-level labels required for training or evaluating patch classifiers.
Performance Metrics
The performance of the models was evaluated using accuracy and the Area Under the Receiver Operating Characteristic Curve (AUCROC).
4.3 Contextual Positive Pair Sampling
Mining neighboring patches based on distance in WSIs carries the risk of sampling false positive pairs, referred to as mismatches. To quantify this risk, we calculate the percentage of mismatches as the sampling distance increases, as illustrated in Figure 2. Note that we calculated mismatching percentages only for WSIs annotated as non-benign (malignant for Camelyon16; Dyplasia and malignant for the private datasets), as benign WSIs contain only benign patches. Additionally, we excluded the TCGA dataset due to the absence of patch-level labels.
Figure 2: Percentage of mismatches as a function of Chebyshev distance for three histopathology datasets (Stomach, Colon, and Breast). TCGA (Lung) was excluded because it does not contain patch-level annotations.
As the distance grows, the likelihood of mismatches rises, reflecting that patches drawn from more distant regions are more likely to belong to different classes or tissue types. Based on this finding, we sampled patches adjacent to the pivot patches as positive pairs, enforcing a maximum Chebyshev distance of 1. During training, neighbor patches are selected randomly. Moreover, we set $\alpha=0.5$ to maintain an equal balance between contextual similarity and transformation invariance. The impact of distance constraints and $\alpha$ was further examined in our ablation studies. Additionally, we explored a scenario where neighbor patches were sampled randomly without any distance constraints.
4.4 Slide-level Classification
Table 1 compares standard positive pair sampling with two contextual variants: Context($\infty$), which samples neighbor patches without distance constraints, and Context($1$), which restricts sampling to patches within a Chebyshev distance of 1. Across all datasets and SSL methods, Context($1$) consistently outperforms the standard approach in both accuracy and AUROC, highlighting the effectiveness of leveraging immediate spatial context.
Table 1: Slide-level performance across datasets, reporting accuracy and AUROC for Positive Pair Sampling strategies. Gray cells indicate contextual approaches: ($\infty$) for unrestricted distance sampling and ($1$) for Chebyshev distance of $1$. BT denotes Barlow Twins. Bold indicates highest value. Results are averaged over five runs.
Dataset SSL Accuracy AUROC
Standard Context($\infty$) Context($1$) Standard Context($\infty$) Context($1$)
Private (Stomach) BT 0.827 0.836 0.893 0.941 0.947 0.966
BYOL 0.813 0.833 0.902 0.954 0.964 0.972
VICReg 0.831 0.816 0.893 0.947 0.945 0.960
DINOv2 0.867 0.893 0.896 0.962 0.967 0.972
Private (Colon) BT 0.820 0.844 0.891 0.926 0.945 0.971
BYOL 0.802 0.756 0.916 0.922 0.930 0.987
VICReg 0.807 0.809 0.916 0.926 0.931 0.981
DINOv2 0.789 0.841 0.867 0.919 0.965 0.954
Camelyon16 (Breast) BT 0.710 0.710 0.763 0.656 0.658 0.722
BYOL 0.757 0.767 0.798 0.751 0.738 0.753
VICReg 0.757 0.716 0.798 0.726 0.705 0.777
DINOv2 0.788 0.817 0.8346 0.769 0.799 0.863
TCGA (Lung) BT 0.773 0.771 0.825 0.839 0.835 0.909
BYOL 0.763 0.760 0.849 0.847 0.844 0.920
VICReg 0.770 0.769 0.857 0.846 0.846 0.914
DINOv2 0.701 0.737 0.833 0.785 0.75 0.893
The performance of Context($\infty$) is more variable—it sometimes matches or exceeds the standard approach but occasionally underperforms, suggesting that unconstrained sampling may introduce noise or less relevant pairs. On the other hand, the magnitude of improvement with Context($1$) varies across datasets. For instance, the Colon dataset sees accuracy gains exceeding 10% (from 0.807 to 0.916 with VICReg), while the Breast dataset improves by about 4% (from 0.757 to 0.798 with VICReg). This variation likely reflects differences in dataset characteristics, such as the presence of multiple tissue samples per WSI or inherent structural complexity. Despite these variations, the consistent advantage of Context($1$) across all datasets underscores the importance of spatially coherent sampling for learning more robust histopathology representations. These findings suggest that leveraging spatial context in SSL not only improves performance in slide-level classification but also provides a biologically intuitive framework for annotation-free learning in digital pathology.
4.5 Linear Probing Evaluation
Table 2: Patch-level performance (Linear Probing) across datasets, reporting accuracy and AUROC for positive pair sampling strategies. Standard refers to self-supervised learning without contextual pairing (i.e., only augmentation-based positive pairs), while Context($1$) refers to our method using spatially adjacent pairs (Chebyshev distance of 1). Supervised indicates fully supervised training; the first value corresponds to a ResNet-18 backbone, and the second (in the DINOv2 row) to a ViT-Tiny backbone. Results are averaged over five runs.
Dataset SSL Accuracy AUROC
Supervised Standard Context($1$) Supervised Standard Context($1$)
Private (Stomach) BT 0.828 0.739 0.783 0.942 0.912 0.928
BYOL 0.774 0.809 0.913 0.938
VICReg 0.803 0.821 0.933 0.944
DINOv2 0.830 0.750 0.762 0.944 0.896 0.904
Private (Colon) BT 0.818 0.840 0.827 0.943 0.939 0.935
BYOL 0.861 0.885 0.950 0.963
VICReg 0.859 0.862 0.950 0.952
DINOv2 0.878 0.797 0.858 0.962 0.921 0.945
Camelyon16 (Breast) BT 0.885 0.764 0.756 0.947 0.831 0.817
BYOL 0.777 0.782 0.834 0.821
VICReg 0.792 0.798 0.860 0.877
DINOv2 0.910 0.797 0.858 0.964 0.921 0.945
To further evaluate the quality of the learned representations, we conducted linear probing using both ResNet-18 and ViT-Tiny backbones. As shown in Table 2, incorporating spatially contextual positive pairs (Context($1$)) consistently improves performance over standard SSL across most datasets and methods.
In the Stomach and Colon datasets, contextual pairing improves AUROC and accuracy in nearly all cases, especially for BYOL and VICReg, which benefit from the spatial coherence in tissue morphology. Notably, BYOL on Colon improves from 0.861 to 0.885 in accuracy and from 0.950 to 0.963 in AUROC. Similar trends are observed in VICReg and DINOv2. For the Camelyon16 (Breast) dataset, improvements are more modest and less consistent. In some cases, such as BT, contextual sampling slightly reduces performance, suggesting that local morphological context may be less discriminative in this dataset, possibly due to higher heterogeneity in slide content or smaller ROI areas.
Across all datasets, standard SSL representations perform worse than both the supervised baseline and our contextual method, reinforcing that augmentation-based invariance alone may be insufficient for modeling histopathology structures. Importantly, contextual SSL nearly closes the gap with fully supervised training and, in some cases (e.g., BYOL and VICReg on Colon), outperforms the supervised ResNet-18 model.
Finally, ViT-Tiny models (DINOv2 rows) exhibit notable improvements with contextual pairing (e.g., from 0.797 to 0.858 accuracy on Breast), indicating that transformer-based encoders can also benefit from spatial coherence, though they remain more sensitive to context quality.
4.6 Ablation Studies
4.6.1 Proportion of Contextual Similarity
Our loss balances contextual similarity and transformation invariance through the mixing parameter $\alpha$. Figure 3 shows that incorporating contextual signals ($\alpha>0$) generally improves accuracy across datasets and methods. The best performance typically occurs at intermediate values ($\alpha=0.25$ or $0.5$), while relying solely on contextual pairs ($\alpha=1.0$) often degrades accuracy, particularly for DINOv2. This suggests that contextual cues alone may lack sufficient diversity, especially for transformer-based models.
Among methods, VICReg shows stable improvements, while BYOL is more sensitive to $\alpha$. BT performs well at moderate $\alpha$ values, and DINOv2 benefits from some contextual information but drops sharply when overused.
Figure 3: Accuracy gain relative to the baseline (without contextual information; i.e., $\alpha=0$) for each SSL method across four datasets. Results are averaged over five runs. Dark blue, orange, green, and purple lines correspond to Barlow Twins (BT), BYOL, VICReg, and DINOv2, respectively.
4.6.2 Effect of Neighboring Distance
Figure 4: Effect of neighboring distance on slide-level classification accuracy measured via accuracy gain (%). Error regions represent the standard deviation.
We investigated how the spatial distance between pivot and neighbor patches affects the performance of self-supervised learning (SSL) in slide-level classification. From a pool of 100,000 pivot patches, we sampled neighbors within defined distance thresholds to capture both local and broader tissue context. As shown in Figure 4, accuracy gains are highest when neighbor patches are sampled from short distances, with performance peaking consistently at distance 1 across all datasets and SSL methods. Beyond this point, accuracy generally declines, particularly at distances greater than 5, likely due to reduced contextual coherence and increased chances of mismatched or noisy pairs. The effect is consistent across Barlow Twins, BYOL, and VICReg, although the extent of gain varies by dataset and method. This variation may reflect differences in tissue structure and intra-slide heterogeneity.
Interestingly, the Stomach and Colon datasets exhibit a partial rebound in accuracy at the largest tested distance (30), possibly due to the presence of multiple tissue fragments per slide, which can create structurally coherent regions even at wider separations. Nonetheless, our results suggest that immediate spatial proximity (i.e., $distance=1$) offers the most reliable and biologically meaningful context for representation learning.
Due to computational constraints, we excluded DINOv2 from these experiments. Exploring its behavior under spatial proximity-based sampling remains promising future work.
5 Discussion
5.1 Spatial Context-Driven Positive Pair Sampling
This work presents a spatial context-driven positive pair sampling strategy tailored for self-supervised learning (SSL) in whole-slide image (WSI) analysis. While conventional SSL methods define positive pairs through synthetic augmentations, such approaches often fail to capture the structural continuity and local semantic coherence of histopathology data. Our method addresses this by explicitly incorporating spatial proximity into the positive pair selection process, leveraging the intrinsic grid structure of WSIs to guide representation learning.
Experimental results from both slide-level classification and patch-level linear probing confirm the effectiveness of our approach. Across datasets and SSL frameworks, incorporating contextual positive pairs, especially those sampled from immediate spatial neighbors (Chebyshev distance of 1), consistently improves downstream performance. For instance, Barlow Twins and VICReg show 4 to 8 percent accuracy gains on slide-level classification tasks in the stomach and colon datasets when using spatially adjacent pairs, compared to standard augmentation-only baselines. These improvements suggest that morphologically similar regions in close spatial proximity provide a more informative training signal than synthetic augmentation alone.
Our method also differs fundamentally from prior work such as SimTriplet, which uses spatial adjacency in histopathology but with a different formulation. SimTriplet employs a triplet architecture and custom loss function to enforce both intra- and inter-sample similarity, requiring a specialized model structure and training setup. In contrast, our method is architecture-agnostic and compatible with a wide range of SSL losses. It integrates directly into joint embedding frameworks such as BYOL, Barlow Twins, VICReg, and DINOv2 by combining augmentation-based and spatially adjacent positive pairs within a unified loss formulation. This design removes the need for handcrafted triplets or predictors, making our approach lightweight, flexible, and easy to adopt.
5.2 Relevance to Pathology Foundation Models
Recent advances in digital pathology have introduced large-scale foundation models that aim to generalize across datasets, tissue types, and diagnostic tasks [21, 26, 28]. These models typically rely on SSL objectives applied at massive scale. While our method is not intended as a replacement for such models, it offers a biologically motivated enhancement to existing pretraining strategies by incorporating spatial coherence into positive pair selection.
By aligning SSL sampling with the inherent architecture of tissue, our approach complements the goals of foundation models, particularly in settings where local morphological consistency is a strong learning signal. Its modular design allows for straightforward integration into pretraining pipelines without requiring architectural changes or significant computational overhead. As foundation models continue to scale, integrating spatial context into their learning objectives may further improve their data efficiency, interpretability, and clinical relevance.
5.3 Limitations and Future Directions
While our method demonstrates consistent performance improvements across datasets and SSL frameworks, several limitations remain. First, all experiments were conducted at a single magnification level using fixed patch sizes. Tissue morphology can exhibit scale-dependent features, particularly in structures with hierarchical organization. Future work should explore how spatial context-driven sampling behaves across multiple resolutions and its sensitivity to patch size and scale.
Second, while we evaluate the method across several SSL frameworks, each framework was tested with a single backbone architecture (e.g., ResNet-18 for Barlow Twins, BYOL, and VICReg; ViT-Tiny for DINOv2). Further testing with deeper or alternative backbones (e.g., ResNet-50, ConvNeXt, Swin Transformer) would provide stronger evidence for the approach’s framework- and architecture-agnostic applicability.
Third, our sampling strategy uses a fixed Chebyshev distance to define spatial neighbors. Although we conduct empirical ablation across multiple distances, a more adaptive approach—one that adjusts sampling radius based on local tissue density, slide size, or training dynamics—could yield more informative and robust positive pairs, particularly in heterogeneous or sparse tissues.
Fourth, due to computational constraints, we excluded DINOv2 from the spatial proximity experiments. Given the strong performance of transformer-based architectures in SSL, it remains an open and valuable direction to investigate how DINOv2 responds to spatial context-driven pair sampling.
Finally, our evaluation focuses primarily on classification tasks using attention-based MIL and linear probing. The potential benefits of spatially informed representations for other downstream tasks (such as clustering, retrieval, or feature interpretability) remain unexplored. Future research should examine how spatial context impacts representation disentanglement, interpretability, and performance in multi-task or clinical applications.
6 Conclusion
We proposed a spatial context-driven positive pair sampling strategy to enhance self-supervised learning (SSL) for whole-slide image (WSI) analysis. By leveraging the natural spatial coherence of tissue morphology, our approach introduces biologically grounded supervision into existing SSL pipelines without modifying their core architectures or objectives. Experimental results demonstrate that incorporating spatially adjacent positive pairs improves downstream performance, particularly in slide-level classification under Multiple Instance Learning (MIL) and patch-level classification via linear probing. Ablation studies show that smaller neighbor distances, especially immediate adjacency, yield the most consistent gains, though the optimal threshold can vary by dataset. Importantly, relying exclusively on spatial context may degrade representation quality, highlighting the value of balancing contextual and augmentation-based signals. Our goal was not to surpass state-of-the-art benchmarks, but rather to introduce a modular, interpretable improvement to SSL that reflects domain-specific structure. The findings suggest that spatial context is a valuable signal for learning more diagnostically relevant features in computational pathology. This strategy is particularly well-suited for large-scale, annotation-limited settings and could complement emerging pathology foundation models. Future work should explore multi-scale integration, adaptive sampling mechanisms, broader tissue diversity, and evaluation across additional tasks such as retrieval, segmentation, and clinical interpretability. Overall, our method offers a scalable and generalizable enhancement to self-supervised learning in histopathology.